OpenStudy (anonymous):

can someone help me with this one... Compare and contrast the two quadratic equations below. describe the following: The direction each parabola opens The vertex of each parabola y = x2 − 2x y = −2x2 + 4x − 3

OpenStudy (campbell_st):

the general form is f(x) = ax^2 + bx + c the 1st equation concave up, since a>0 f(x) = x(x - 2) has x intercepts at x = 0 and x = 2 the line of symmetry is x = 1 y intercept of y = 0 and a minimum value of y = -1 the 2nd equation concave down since a<0 is negative definate since the discriminant \[\Delta = 4^2 - 4\times(-2)\times(-3) = -10\] which means the curve doesn't cut the x - axis has a line of symmetry of x = 1 has a maximum value of y = -1 y intercept at y = -3 hope this helps

OpenStudy (anonymous):

sure thanks <3